OpenStudy (anonymous):

Translate into an equation and show steps to solve. Money is invested in a "guaranteed fund" at 12% simple interest. After one year, there is $840 in the account. How much was originally invested?

OpenStudy (kropot72):

Let the amount originally invested = A 840 = A + (0.12 * A) Can you solve this equation to find A ?

OpenStudy (anonymous):

is 3500 the correct answer?

OpenStudy (kropot72):

A * 1.12 = 840 \[A=\frac{840}{1.12}=?\]

OpenStudy (anonymous):

james, if originally u saved 3500 and after a year u got 840, that is not make sense

OpenStudy (anonymous):

o0h thats make sense if u went to ATM much :D

OpenStudy (kropot72):

@JamesR4494 Can you calculate\[\frac{840}{1.12}\] to find A ?

OpenStudy (anonymous):

750, How'd you get 1.12 though?

OpenStudy (anonymous):

(1+r)^-t

OpenStudy (kropot72):

1.12 is obtained from the equation 840 = A + (0.12 * A) = A(1 + 0.12) = A * 1.12

OpenStudy (anonymous):

I still don't understand :/

OpenStudy (anonymous):

do u know the formula?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

the original money + interest = the money in the end, do u understand?

OpenStudy (anonymous):

So we have A + 0.12 = $840

OpenStudy (kropot72):

A + 0.12A = 840

OpenStudy (anonymous):

yes, but the formula for interest is (1+ 12%)^-1 year